OpenStudy (anonymous):

Find the slope of the line with these given points: (0, 4) and (2, 5). -1/2 1/2 -2 2

OpenStudy (anonymous):

slope of a line, m= (y2-y1)/(x2-x1) Thus take x1 = 0, y1 = 4, x2 = 2 and y2 = 5 Put the values in m, you will get slope,m= 1/2